OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9(F - 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is.

OpenStudy (anonymous):

1) C = (5/9)(F-32) F - 32 = (9/5)C F = (9/5)(C) + 32 2) Plug and solve: The answer is 68. Try for yourself.

OpenStudy (lilg132):

9/5 = 1.8 F = 9/5 * 20 + 32 or F = 1.8 * 20 + 32